What would the answer be for 12+2(k+3)=3k-6?
Could you please show the work so i can understand it.

Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(48568)

12+2(k+3)=3k-6
Use the distributive property to get:
12+2k+6 = 3k-6
Combine like terms on the left side to get:
2x+18=3k-6
Add 6 to both sides to get:
2x+24=3x
Subtract 2x from both sides to get:
x=24
Cheers,
Stan H.

Answer by Edwin McCravy(6939)

What would the answer be for

                        12 + 2(k + 3) = 3k - 6

First get rid of the parentheses on the left by 
multiplying the +2 times both terms inside it, 
the k and the +3, giving +2k + 6

                          12 + 2k + 6 = 3k - 6

Add the 12 and the +6, getting 18 

                              18 + 2k = 3k - 6

Get rid of the 18 on the left by adding -18 to both 
sides:


                              18 + 2k = 3k -  6
                             -18            -18
                             ------------------
                                   2k = 3k - 24

Get rid of the 3k on the right by adding -3k to both 
sides:

                                   2k = 3k - 24
                                  -3k  -3k  
                                  -------------
                                   -k =     -24

Now write -k as -1k

                                  -1k = -24

Now we get rid of the -1 coefficient of k, we divide
both sides by -1:

                                 -1k     -24
                                ————— = —-———
                                 -1       -1

The -1's cancel on the left, leaving only k.
-24 divided by -1 on the right gives +24 on the
right and we are left with:

                                    k = 24

The solution is 24. 

Edwin